### v1

# Best Agent Identification for General Game Playing

Abstract
We present an efficient and generalised procedure to accurately identify the best performing algorithm for each sub-task in a multi-problem domain. Our approach treats this as a set of best arm identification problems for multi-armed bandits, where each bandit corresponds to a specific task and each arm corresponds to a specific algorithm or agent. We propose an optimistic selection process based on the Wilson score interval (Optimistic-WS) that ranks each arm across all bandits in terms of their potential regret reduction. We evaluate the performance of Optimistic-WS on two of the most popular general game domains, the General Video Game AI (GVGAI) framework and the Ludii general game playing system, with the goal of identifying the highest performing agent for each game within a limited number of trials. Compared to previous best arm identification algorithms for multi-armed bandits, our results demonstrate a substantial performance improvement in terms of average simple regret. This novel approach can be used to significantly improve the quality and accuracy of agent evaluation procedures for general game frameworks, as well as other multi-task domains with high algorithm runtimes.
Keywords: General game playing, best arm identification, multi-armed bandits, Wilson score, agent evaluation
1 Introduction
The field of general game playing focuses on the development of AI agents that are able to effectively play a wide variety of games (Genesereth et al., 2005). This can often be viewed as a multi-problem domain, where each agent needs to perform well across a collection of individual tasks. Designing artificial general agents that are able to successfully play a wide range of games has long been a core research area for artificial intelligence and machine learning, leading to significant breakthroughs in the areas of deep reinforcement learning (Torrado et al., 2018; Goldwaser and Thielscher, 2020), search and planning (de Waard et al., 2016; Frydenberg et al., 2015), transfer learning (Banerjee and Stone, 2007; Soemers et al., 2023), evolutionary algorithms (Gaina et al., 2017), among many others. However, as the number of games and agents developed for these frameworks increases, so too does the evaluation time needed to reliably assess the performance of each agent. Many agents do not play identically when repeatedly attempting a game, often requiring a large number of trials to obtain an accurate performance estimate (Bontrager et al., 2021). While there is often “no free lunch” when it comes to producing more results in less time (Ashlock et al., 2017), several approaches have been proposed to reduce the number of unnecessary evaluations using prior domain knowledge (Stephenson et al., 2020; Aitchison et al., 2023).
Rather than needing to accurately assess the game playing abilities of each agent across all games, our aim is to develop a domain agnostic approach that can efficiently identify the best performing algorithm (i.e., agent) for each task (i.e., game) within a defined set. Knowing the best agent for each game can be useful for several purposes, such as training prediction models for portfolio/ensemble agents or detecting skill gaps in the existing agent suite (Anderson et al., 2019; Stephenson and Renz, 2017). We frame this challenge as a multi-bandit best arm identification problem (Gabillon et al., 2011), where each bandit represents a specific game and each arm represents an available agent for playing that game.
In this paper, we propose a novel multi-bandit best arm identification algorithm called Optimistic-WS. This approach combines an “optimism in the face of uncertainty” philosophy with a generalised interpretation of the Wilson score interval, selecting arms that have the highest potential for improving our best arm estimation (i.e., minimising simple regret) across all bandits. We evaluate and compare the performance of our Optimistic-WS approach against prior multi-bandit best arm identification algorithms for two of the most popular general game domains. This includes the General Video Game AI (GVGAI) framework (Perez-Liebana et al., 2019) and the Ludii general game playing system (Piette et al., 2020), both of which are frequently used within the academic community for developing and comparing general game playing techniques (Perez-Liebana et al., 2016; Piette et al., 2023). Our results demonstrate that our proposed Optimistic-WS approach provides significant performance improvements with regards to best agent identification when compared to alternative state-of-the art techniques.
The remainder of this paper is organised as follows: Section 2 formalises the multi-bandit best arm identification problem and its applicability to best agent identification for general game playing. Section 3 describes relevant prior work on general game systems, efficient agent evaluation procedures, best arm identification algorithms, and the Wilson score interval. Section 4 defines our proposed Optimistic-WS approach. Section 5 provides experiment details and performance evaluations using the GVGAI and Ludii general game playing domains. Section 6 summarises our results, highlighting potential areas for further investigation and improvement.
2 Problem Setup
In this section, we provide a formalisation of the multi-bandit best arm identification problem and its applicability to our intended domain of general game playing.
2.1 Multi-Bandit Best Arm Identification
Best-arm identification is a variation of the more traditional stochastic multi-armed bandit problem. Rather than maximising the cumulative reward over multiple arm pulls, the aim of best-arm identification is to maximise the reward of a single arm pull after a defined exploration period (i.e., select the arm with the highest expected reward after a limited number of prior arm pulls). This problem can be expanded further into the multi-bandit best arm identification problem, where instead of a single bandit machine there are now multiple bandits available to choose from, each with their own selection of arms and reward distributions. Rather than identifying the single best arm, the goal of multi-bandit best arm identification is to identify the best performing arm for each individual bandit. A more formalised definition of the multi-bandit best arm identification problem is provided below, with the majority of terms borrowed from (Gabillon et al., 2011).
The setup for a multi-bandit environment can be defined as follows. Let $M$ be the number of bandits and $K$ represents the number of arms available for each bandit. Each individual bandit-arm pair $(m,k)$ can be characterised by a stationary reward distribution $\nu_{mk}$, bounded in $[0,b]$ and with mean reward $\mu_{mk}$. Assuming that each bandit has a unique best arm, $k_{m}^{\star}$ and $\mu_{m}^{\star}$ represent the index and mean of the best arm for bandit $m$ (i.e., $k_{m}^{\star}=\operatorname*{argmax}_{1\leq k\leq K}\mu_{mk},\ \mu_{m}^{\star}=\max_{1\leq k\leq K}\mu_{mk}$).
The problem of Best Arm Identification occurs within a multi-bandit environment when the true reward distributions $\{\nu_{mk}\}$ of our bandits are unknown, and is often framed as a prediction task. During each round $t=1\dots n$, we can pull a single bandit-arm pair $I(t)=(m,k)$ and observe an independent sample reward drawn from the distribution $\nu_{I(t)}$. Let $T_{mk}(t)$ be the number of times that bandit-arm pair $(m,k)$ has been pulled by the end of round $t$, then the mean reward of this bandit-arm pair can be estimated as $\hat{\mu}_{mk}(t)=\frac{1}{T_{mk}(t)}\sum_{s=1}^{T_{mk}(t)}X_{mk}(s)$, where $X_{mk}(s)$ is the $s$-th sample observed from $\nu_{mk}$. After all $n$ rounds have finished, for each bandit $m$ we return the arm $J_{m}(n)=\operatorname*{argmax}_{k}\hat{\mu}_{mk}(n)$ with the highest estimated mean reward.
To evaluate the performance of our approach after $n$ rounds, for each bandit we compare the expected reward for our selected arm against that of the true best arm. This comparison is usually done by calculating the average difference in expected rewards for each arm pair, known as the simple regret $r(n)=\frac{1}{M}\sum_{m=1}^{M}(\mu_{m}^{\star}-\mu_{mJ_{m}(n)})$. In certain situations, the percentage of incorrectly identified best arms may be a more desirable measure of performance, $e(n)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{P}(J_{m}(n)\neq k_{m}^{\star})$. For our case study, we will focus on using the average simple regret $r(n)$ as our measure of algorithm performance.
2.2 Best Agent Identification for General Game Playing
The above problem of multi-bandit best arm identification is analogous to determining the best performing algorithms across a variety of different tasks. Each task can be treated as a single multi-armed bandit, with each arm representing a potential algorithm that can perform this task (albeit with varying degrees of success). Using this framing allows us to represent our task of identifying the best agent for each game within a general game framework as a multi-bandit best arm identification problem. Each game is treated as its own multi-arm bandit, with each arm of this bandit representing a different game playing agent. A pull of a specific arm for a specific bandit corresponds to running a single trial for the associated game-agent pair, with the agent’s outcome at the end of this trial determining the reward obtained.
2.2.1 Domain Considerations
While prior approaches for solving multi-bandit best arm identification problems have been proposed for the general case described above, our specific focus on general game playing introduces several additional considerations:
Reward Bounds - One assumption we can make for the domain of general game playing is that outcome for an agent at the end of a trial is always between 0 and 1 (inclusive). Typically an outcome of 0 indicates a loss and an outcome of 1 indicates a win, with other outcomes between these indicating either a draw or positional ranking among multiple players. This means that the value of $b$ in our reward distribution bounds will always be equal to 1 (i.e., all rewards bounded in $[0,1]$).
Unequal Number of Arms - Another consideration is that the above definition of a multi-bandit environment assumes an equal number of arms $K$ for each bandit. However, in our application to general game playing it is often the case that certain agents may not be able to play every game available. In these situations, we can simply apply a generalisation to the above definitions that replaces the static value $K$ with a bandit-specific value $K_{m}$ that represents the number of arms available for a given bandit $m$.
Multiple Best Arms - Lastly, when comparing multiple agents for a single game, the assumption that there will always be a unique best arm $k_{m}^{\star}$ for any bandit $m$ is unlikely to hold. For example, any game where several agents are always able to win will result in an equally high reward estimation of 1.0 for multiple arms. To account for this, we define our $\operatorname*{argmax}$ function to always return a single best argument combination, chosen at random from the highest value arguments. We will also need to modify the average probability of error measure $e(n)$, so that selecting any arm with the highest expected reward $\mu_{m}^{\star}$ will be recorded as a correct prediction (i.e., $e(n)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{P}(\mu_{m}^{\star}\neq\mu_{mJ_{m}(n)})$).
3 Background
3.1 General Game Systems
While many different general game domains have been proposed over the past decade, arguably the two most popular for academic research have been the General Video Game AI framework (GVGAI) and the Ludii general game playing system.
3.1.1 General Video Game AI Framework
The General Video Game AI (GVGAI) framework contains over 100 simple arcade-style video games, described using the Video Game Description Language (VGDL) (Perez-Liebana et al., 2019). The GVGAI framework has been the subject of a long running general game playing competition (Perez-Liebana et al., 2016), leading to several dozen agents being developed for it over the years (Pérez-Liébana et al., 2016; Soemers et al., 2016; Gaina et al., 2017; Weinstein and Littman, 2012; Mendes et al., 2016). While there are several auxiliary research tracks available for the GVGAI framework, in this paper we will focus solely on the games and agents developed for the main single-player planning track. Each game includes five distinct levels, which all use the same base set of mechanics but vary the number and initial locations of objects. Previous analysis on agent performance across the corpus of GVGAI games has observed that agents often produce highly varied reward distributions (Bontrager et al., 2021). Many of the games available within the GVGAI framework also have some form of stochasticity, such as randomised items or enemy behaviour, making agent evaluation an inherently noisy process. These factors, along with the time consuming nature of running dozens of trials across thousands of game-agent pairs, makes the GVGAI framework an ideal case study for best agent identification.
3.1.2 Ludii general game playing system
In contrast to the GVGAI framework’s focus on video games, the Ludii general game playing system instead provides over 1000 board and puzzle games (Piette et al., 2020). This makes Ludii one of the largest collections of playable games for AI research purposes. Games within Ludii are described using the ludemic game description language, which supports a wide range of mechanics and game types. This includes stochastic or hidden information, alternating or simultaneous moves, and player counts ranging from 1 to 16 (Stephenson et al., 2019). Ludii has also been used as a competition framework for promoting the development of general game playing agents (Piette et al., 2023), as well as research into general game heuristics (Stephenson et al., 2021). Much like the GVGAI framework, agent evaluation in Ludii is a lengthy and imperfect process, often requiring maximum turn limits to even guarantee that a game will end. These aspects make Ludii a prime candidate for applying more efficient best agent identification algorithms.
3.2 Efficient Agent Evaluation
While we believe that this is the first time the task of best agent identification for general game systems has been framed from a multi-arm bandit perspective, this section discusses alternative techniques for more efficiently evaluating general game agents.
3.2.1 Game Subsets
Rather than evaluating all agents across the full suite of available games, one alternative is to identify a representative subset of games to evaluate on. This approach has been previously investigated for both the General Video Game AI framework (Stephenson et al., 2020) and Arcade Learning Environment (Aitchison et al., 2023). This technique has also been mentioned as possible future work for the Ludii general game playing system (Piette et al., 2021; Stephenson et al., 2023) but has yet to be properly investigated. While this approach has a similar motivation to our current problem, being that testing all game-agent pairs a sufficient number of times to get a reliable indication of performance is computationally expensive, our goal is to instead focus on identifying the best agent for each individual game, rather than assessing the general performance of each agent. These approaches also use domain-specific measures of game and/or agent similarity to identify a representative game subset, whereas our presented approach requires no such domain knowledge.
3.2.2 Active Learning
Active Learning is a branch of Machine Learning where labelled training instances are selected through an intelligent process (Settles, 2011). This typically occurs when there is a large quantity of unlabelled instances, but labelling any given instance is costly to perform. An active learning algorithm attempts to identify which instances should be labelled to provide maximal training benefit. This situation is highly similar to our described problem, but instead of an abundance of unlabelled instances that need labelling we instead have a noisy and/or probabilistic labelling process. In both cases, the core motivation that obtaining a large number of accurate instance labels is prohibitively expensive remains the same. While previous research has demonstrated that active learning approaches can be used to minimise reward estimate uncertainty for multi-armed bandits (Deng et al., 2011), the problem of best arm identification has not yet been explored.
3.3 Best Arm Identification
Several prior algorithms for the problem of best arm identification have been presented over the past few decades, either focusing on the problem from a general perspective or for a specific set of circumstances. Common variations include correlated arm rewards (Gupta et al., 2021; Kazerouni and Wein, 2021; Hoffman et al., 2014), identifying the best-$N$ arms (Bubeck et al., 2013), identifying the best arm across multiple overlapping groups (Scarlett et al., 2019), and restless arms with evolving states (Narayana Prasad et al., 2022). The majority of the these approaches focus on best arm identification for a single bandit, but can also be generalised to a multi-bandit domain. Out of these past approaches, four algorithms (and variants of them) stand out as being the most applicable to our use case.
3.3.1 GapE
The Gap-based Exploration (GapE) algorithm (Gabillon et al., 2011) is one of the first approaches designed specifically for the multi-bandit best arm identification problem. This algorithm favours arms with an expected reward close to the best performing arm (i.e., those with the smallest “gap”). The authors of this approach also define a variant called GapE-V, which additionally considers the variance in rewards for each arm.
While on the surface this approach seems well suited for our intended application, the specific characteristics of our general game playing domains highlight some potential issues. Prior observations and experience with these general game frameworks has demonstrated that, while there are certain games where agents will frequently achieve different outcomes between repeated trials, a significant portion of games also have a very low variability in agent performance. In these cases, there are often several agents that either always win or always lose the game (i.e., reward distributions for game-agent pairs are asymmetric and centred towards the extremities). This means that there will be multiple arms for a given bandit with a near-zero gap between them and the best performing arm.
For our proposed domain, we care primarily about minimising the average simple regret $r(n)$ of our selected arms. As such, if there are several best or near-best arms for a given game, then the decision of which one we select will have a negligible impact. Because GapE has been designed primarily for minimising the average probability of error $e(n)$, the approach instead focuses a large number of arm pulls on these equally high performing arms (i.e., the algorithm attempts to identify the one true “best” arm, rather than being directly concerned with reducing the simple regret of the bandit). This difference often leads to a sub-optimal arm selection strategy from GapE when attempting to minimise the average simple regret across all bandits.
3.3.2 UCB-E
The UCB-E algorihtm (Audibert et al., 2010) is a highly exploring variant of the Upper Confidence Bound (UCB) selection policy (Auer, 2003). One of the most common algorithms for standard (i.e., cumulative reward) multi-arm bandit problems is the UCB1 algorithm (Auer et al., 2002), which is able to effectively balance exploration and exploitation. However, best arm identification can be considered a pure exploration problem, as the rewards obtained from each arm during runtime do not affect our performance score (only the final arm pull does). UCB-E attempts to increase the degree of exploration performed by UCB, based on a defined exploration value $a$. If $a=2\cdot\log(n)$, where $n$ is the current number of rounds, then UCB-E is equivalent to UCB1, with the value for $a$ being increased to incentivise more exploration. The effectiveness of UCB-E is highly dependant on the value chosen for $a$, for which the optimal value is often unknown and cannot be easily estimated from past observations. While originally designed for best arm identification in a single bandit case, UCB-E can be generalised to the multi-bandit environment by selecting bandits uniformly and then pulling arms within each bandit using the UCB-E algorithm (Gabillon et al., 2011).
3.3.3 Successive Rejects
The Successive Rejects algorithm (Audibert et al., 2010) repeatedly removes the worst performing arm over multiple rounds of selection until just a single arm remains, a process commonly referred to as “action elimination”(Jamieson and Nowak, 2014). However, in order for this and other action elimination algorithms to be used to their fullest effectiveness, the total number of arm pulls must be known in advance. This lack of a “stop anytime” functionality is a significant downside compared to previous approaches, particularly for our intended use case of general game playing. The amount of time needed to complete a single game trial can be highly variable, meaning that we cannot easily estimate the exact number of trials that can be run within an allotted time. The generalisation of this approach to multiple bandits is also non-optimal, requiring the number of rounds (i.e., arm pulls) to be equally split among all bandits at the start of the process, rather than to bandits with the highest potential benefit during runtime.
3.3.4 Sequential Halving
The original Sequential Halving algorithm (Karnin et al., 2013) works in a manner similar to that of the Successive Rejects algorithm, where the worst performing arms of each bandit are removed at the end of each of multiple rounds. However, rather than removing a single arm at the end of each round, half of the remaining arms are removed instead. Sequential Halving also suffers from the same core limitation as Successive Rejects, namely being the requirement that the number of available arm pulls must be known in advance.
Anytime Sequential Halving (Sagers et al., 2025) provides a variant approach that avoids this limitation. This variant performs an initial run of the Sequential Halving algorithm, with a limited budget of arm pulls equal the minimum amount required for a single complete run. Once complete, this “minimal pull run” process is repeated, with subsequent reward estimates for each arm incorporating all prior pull results from previous runs. While this Anytime Sequential Halving variant offers the desired stop anytime functionality, it still suffers from being unable to generalise effectively to the multi-bandit domain. Similar to Successive Rejects, arm pulls must be divided evenly across all bandits, resulting in an equal number of trials for each game.
3.4 Wilson Score
The Wilson score interval is a binomial proportion confidence interval for estimating the probability of success based on prior Bernoulli (aka. binomial) trial outcomes (Wilson, 1927). The Wilson score interval has previously been shown to perform particularly well compared to other confidence estimates when the number of samples is small, or when the probability of success is close to the extremes (i.e., 0 or 1) (Agresti and Coull, 1998; Wallis, 2013). The Wilson score interval for a specified confidence level $(1-\alpha)$ is defined in formula (1), additionally parameterised by the sample probability of success $\hat{p}$ and number of samples $n$. Note, for the sake of our later algorithms, we assume that if the value for $n$ is zero then the returned Wilson score interval is always $(0,1)$.
$Wilson(\hat{p},n,\alpha)\equiv(w^{-},w^{+})=\frac{\hat{p}+\frac{z^{2}_{\alpha/2}}{2n}\pm z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}_{\alpha/2}}{4n^{2}}}}{1+\frac{z^{2}_{\alpha/2}}{n}}$ (1)
While the Wilson score interval was originally intended only for Bernoulli trials with an outcome of either 0 or 1, it can still be applied in practice to non-Bernoulli trials as long as the outcome is always within the required $[0,1]$ bounds. In this case, the value $\hat{p}$ is replaced with the average result of all trials. The fact that Wilson score intervals are known to work particularly well with highly skewed distributions (Wallis, 2013), suggests that this generalised interpretation may provide a suitable confidence estimate for agent win-rate in our general game playing domain.
4 Optimistic-WS
Based on the considerations and restrictions of our general game playing domain, we define the following functionality and design requirements for our proposed approach:
  • Stop Anytime - We can return the current best identified arm $J_{m}(t)$ for each bandit $m$ at the end of any round $t$. This allows us set the total number of rounds $n$ to an arbitrarily high value and return the value for $J_{m}(t)$ whenever needed, or to increase the total number of rounds and continue improving our predictions after the previous best arm predictions $J_{m}(n)$ have been returned.
  • New Games/Agents - We can add new bandits or arms to our environment during runtime. This should be possible at the start of a round by initialising any required parameters for each new game-agent pair.
  • Reward Bounds - Our approach needs to operate within the required [0,1] reward bounds of our domain.
  • Unequal Number of Arms - Our approach needs to function for sets of bandits with different numbers of available arms $K_{m}$.
  • Simple Regret - Our approach should be focused on minimising the average simple regret measure $r(n)$ of our predicted best arms $J_{m}(n)$.
Based on these criteria, we propose our Optimistic-WS algorithm to tackle the best agent identification problem for general game playing. Algorithm 1 provides the pseudocode for this approach, utilising previously defined variables and terminology. A more descriptive explanation of Optimistic-WS operations and the motivation behind its design is provided below.
Algorithm 1 Pseudocode of the Optimistic-WS algorithm
number of rounds $n$, exploration value $c$
$T_{mk}(0)=0,\hat{\mu}_{mk}(0)=0$ for all bandit-arm pairs $(m,k)$
for $t=1,2,...,n$ do
     Compute $\alpha=K_{m}/(t\cdot c$)
     for each bandit $m$ do
         Compute $\hat{k}_{m}^{\star}=\operatorname*{argmax}_{k\in\{1\dots K_{m}\}}\hat{\mu}_{mk}(t-1)$
         Compute $\hat{\mu}_{m}^{\star}=\max_{k\in\{1\dots K_{m}\}}\hat{\mu}_{mk}(t-1)$
         for each arm $k$ of bandit $m$ do
              Compute $(w^{-}_{mk},w^{+}_{mk})=Wilson(\hat{\mu}_{mk}(t-1),T_{mk}(t-1),\alpha)$
              if $\hat{\mu}_{mk}(t-1)=\hat{\mu}_{m}^{\star}$ then
                  $\Delta_{mk}=\hat{\mu}_{m}^{\star}-w^{-}_{mk}$
              else
                  $\Delta_{mk}=w^{+}_{mk}-\hat{\mu}_{m}^{\star}$
              end if
         end for
     end for
     Draw $I(t)\in\operatorname*{argmax}_{m,k}\Delta_{mk}$
     Observe $X_{I(t)}(T_{I(t)}(t-1)+1)\sim\nu_{I(t)}$
     Update $\hat{\mu}_{I(t)}(t)\leftarrow\hat{\mu}_{I(t)}(t-1)+(X_{I(t)}(T_{I(t)}(t-1)+1)-\hat{\mu}_{I(t)}(t-1))/(T_{I(t)}(t-1)+1)$
     Update $T_{I(t)}(t)\leftarrow T_{I(t)}(t-1)+1$
end for
Return $J_{m}(n)=\operatorname*{argmax}_{k}\hat{\mu}_{mk},\ \forall_{m}\in\{1\dots M\}$
The general philosophy behind our proposed Optimistic-WS algorithm is to take the same “optimism in the face of uncertainty” perspective used for most upper confidence bound selection policies. In this case, we selected the Wilson score interval to estimate the upper and lower confidence bounds $[w^{-}_{mk},w^{+}_{mk}]$ of the expected reward of each bandit-arm pair $(m,k)$. The confidence bounds for a given arm are determined based on its expected return $\hat{\mu}_{mk}$, as well as the number of times it has previously been pulled $T_{mk}$ and the confidence value $\alpha$. Higher values for $T_{mk}$ and $\alpha$ will shrink the confidence interval (vice versa). Due to the fact that $Wilson(\hat{p},n,\alpha)$ always returns the maximal confidence interval $(0,1)$ when $n$ is zero, any arm that has not yet been pulled at least once (i.e., $T_{mk}=0$) will always be the preferred selection choice.
While a fixed confidence value ($\alpha$) could be used for all rounds, we instead chose to apply an adaptive confidence value that decreases as the round number ($t$) increases. This reduction ensures that our arm selection process will continue to explore across all bandits even for a very high number of rounds, and guarantees that our proposed Optimistic-WS algorithm will converge towards the optimal result as the number of rounds increases. Without this, it is possible that the gap between the confidence values of the best and second best arms in a bandit becomes large enough that no arms of this bandit are ever pulled again. In practice, the rate at which our confidence value decreases can be adjusted using a specific exploration value ($c$), with the total number of arms $K_{m}$ used as a stabilising value across different problem domains.
The potential regret change $\Delta_{mk}$ for each bandit-arm pair $(m,k)$ is then calculated by comparing its Wilson score confidence bounds $[w^{-}_{mk},w^{+}_{mk}]$ against the expected reward for the current best predicted arm $\hat{\mu}_{m}^{\star}$. If an arm $k$ is not the current best predicted arm $\hat{k}_{m}^{\star}$ for a bandit $m$, then its potential regret change $\Delta_{mk}$ is equal to the difference between its upper confidence bound $w^{+}_{mk}$ and the expected reward of the best performing arm $\hat{\mu}_{m}^{\star}$. This represents the potential increase in regret that would be obtained by selecting arm $k$ for bandit $m$, if its true expected reward were equal to its upper confidence bound $w^{+}_{mk}$. Conversely, if the arm $k$ is the best predicted arm $\hat{k}_{m}^{\star}$ for a bandit $m$, then its potential regret change $\Delta_{mk}$ is equal to the difference between the expected reward of the best performing arm $\hat{\mu}_{m}^{\star}$ (which in this case will be identical to its own expected reward $\hat{\mu}_{mk}$) and its lower confidence bound $w^{-}_{mk}$. This represents the potential decrease in regret that would be obtained by selecting arm $k$ for bandit $m$, if its true expected reward was equal to its lower confidence bound $w^{-}_{mk}$. In summary, we apply an optimistic perspective on all non-best arms by assuming their upper confidence bound is true, and a pessimistic perspective on all best arms by assuming their lower confidence bound is true.
The bandit-arm pair with the largest $\Delta_{mk}$ value is then pulled to obtain our new reward sample $X_{mk}$. This reward sample is used to update the arm’s expected reward estimate $\hat{\mu}_{mk}$ and increment its pulls counter $T_{mk}$. Once this process has been completed the round is over, and the next round can begin. The current best predicted arm for any bandit $J_{m}(t)$ can be returned at the end of any round $t$, providing an intermediate measure of algorithm performance. New bandits or arms can also be added at the start of any round by updating the values for $M$ and $K_{m}$, as well as initialising all values for $T_{mk}(0\dots t)$ and $\hat{\mu}_{mk}(0\dots t)$ to 0.
4.1 Hyperparameters
4.1.1 Number of rounds
One advantage of our proposed algorithm over alternatives such as Successive Reject, is that the total number of rounds $n$ has no impact on which arms are pulled each round. This means that the predicted best arm for each bandit $J_{m}$ can be returned after any number of rounds, without a loss in prediction performance. This “stop anytime” nature also allows for alternative stopping criteria, such as algorithm runtime or prediction confidence, to be easily enforced. It is also possible to add additional rounds (i.e., increase the value of $n$) from an arbitrary stopping point, again without any loss in performance.
4.1.2 Alpha value
The alpha value used for our Wilson score interval determines the desired level of confidence, where a lower alpha value typically equates with a higher degree of exploration (vice versa). In line with the typical exploration-exploitation trade-off, higher alpha values often perform better in the short term and lower alpha values perform better in the long term.
5 Experiments
5.1 Game-Agent Datasets
While the majority of prior research on best arm identification has focused on proving certain theoretical results for specific algorithms, we will instead demonstrate the practical application and empirical results of our approach on real-world data. Specifically, we will use results obtained from the GVGAI and Ludii domains across a variety of games and agents.
Results for GVGAI were sourced from the work of Stephenson et al. (2020), providing a total of 3,990,760 trials across 108 games and 27 agents. These trials are also split amongst the five distinct levels available for each game, allowing us to include agent results from a subset of levels if desired. To evaluate whether agent performance differences between levels affects our approach’s effectiveness, we define both a GVGAI (all levels) dataset that includes trials for each game across all five levels, as well as a GVGAI (level 1) dataset that includes only the 798,516 trials from the first level of each game.
Results for Ludii were obtained by running 29 variants of an Alpha Beta agent, each using a different heuristic state evaluation function (Stephenson et al., 2021). The Alpha Beta agent was set to Player 1, while any other players were controlled by a standard Upper Confidence bounds applied to Trees (UCT) algorithm (Kocsis and Szepesvári, 2006). Each agent was given 0.5 seconds of thinking time per move. A maximum turn limit of 500 was applied to all games, and exceeding this value results in a draw for all remaining active players. Each agent (i.e., Alpha Beta heuristic) was run for 100 trials on each of the 1085 games included in Ludii v1.3.9, giving a total of 3,434,627 trials. Results were generated using a single general node (AMD EPYC 7551 @2.55Ghz, 256GB DDR4 @2666Mhz) of the DeepThought high-performance computer (Flinders University, 2021).
For our presented experiments, rather than running new trials for a chosen game-agent pair we will instead sample rewards randomly from these datasets of previously performed trials. The number of pre-calculated results included in these datasets is large enough to act as a proxy for the true reward distribution of each game-agent pair, and allows for both an increased number of repeated comparison runs and easily reproducible results. Fixing our result datasets also provides us with a defined “ground truth” measure of agent performance, which is necessary for the simple regret calculations that will be used to evaluate and compare different best-agent identification algorithms.
All source code, game-agent datasets and regret scores used to produce the presented results is publicly available online. [^1] https://github.com/stepmat/Best_Agent_Identification_GGP
Algorithm Summary Hyperparameters
Random Random arm selection across all bandits
Uniform Equal arm selection across all bandits
GapE Selects arms that are close to the current best arm in each bandit Exploration value
$a$ = [1, 2, 4, 8, 16]
GapE-V Alternative version of GapE that also considers reward variance Exploration value
$a$ = [1, 2, 4, 8, 16]
UCB-E A highly exploring version of UCB that selects arms with the highest upper confidence bound ($a=2\cdot\log(n)$ is equivalent to UCB1) Exploration value
$a$ = [2, 4, 8, 16] $\cdot$ $\log(n)$, where $n$ = round number
Successive Rejects* Repeatedly dismisses the worst performing arm in each bandit
Sequential Halving* Repeatedly dismisses half the number of arms in each bandit
Anytime Sequential Halving Repeatedly performs a shortened version of sequential halving, using the minimal number of arm pulls required
Optimistic-WS Selects arms with the highest regret change potential, based on Wilson score estimates Exploration value
$c$ = [1, 2, 4, 8, 16]
*Requires a fixed total number of arm pulls.
Table 1: Best Arm Identification algorithms to compare.
5.2 Algorithm Comparison
Using the GVGAI and Ludii game-agent results datasets described above, we evaluate the performance of the multi-bandit best arm identification algorithms described in Table 1. This includes our proposed Optimistic-WS algorithm, along with all previous alternatives described in Section 3.3. To provide a more standard performance baseline, we also evaluate a Random sampling approach that selects a bandit-arm pair at random each time with equal probability (i.e., with replacement), and a Uniform sampling approach that selects each bandit-arm pair an equal number of times (i.e., without replacement). For algorithms with adjustable hyperparameters, a range of suitable values were tested. Each algorithm and/or hyperparameter value was run on both datasets 10 times, with 50,000 rounds (i.e., arm pulls) in each run. The simple regret of each algorithm was calculated every 1000 pulls, providing us with an average measure of performance across a varying number of rounds.
Given that the optimal initial strategy at the start of any best-arm identification process is to first sample all bandit-arm pairs at least once, we decided to provide a single consistent result from each game-agent pair as an initial win-rate estimate at the start of our experiments. Each best-arm identification approach had its performance estimate for each arm initialised with this first result, as any simple regret measures taken before this time would essentially be identical for all algorithms. The 50,000 arm pull budget provided to each algorithm in these experiments is taken in addition to this initial single pull result for each arm.
To prevent cluttering in our presented results, we have selected only the overall best performing hyperparameter value for each algorithm to compare. More specifically, for GapE ($a$ = 2), for GapE-V ($a$ = 1), for UCB-E ($a$ = 2), and for Optimistic-WS ($a$ = 16). The decision of which hyperparameter value was the “best performing” was determined by whichever value had the lowest average simple regret for the highest number of rounds. In practice, this choice of hyperparameter value had little impact on the effectiveness of our Optimistic-WS algorithm, which always managed to outperform all other approaches. Full results for all hyperparameter values are provided in the Appendix.
5.3 Results
5.3.1 GVGAI
The average regret of each best agent identification algorithm for both the GVGAI (all levels) and GVGAI (level 1) datasets are shown in Figures 1 and 2 respectively. From these results we can see that our proposed Optimistic-WS algorithm provides a substantial performance improvement compared to all previous approaches for any number of rounds. When applied to the GVGAI (all levels) dataset, the simple regret of our Optimistic-WS approach is approximately 35.5% smaller than that of the second best approach (UCB-E), and 67.9% smaller than the more typical uniform sampling approach (averaged across the full 50,000 rounds). The improvement provided by Optimistic-WS is even more pronounced for the GVGAI (level 1) dataset, with our simple regret being 47.0% smaller than UCB-E and 70.4% smaller than uniform sampling over the same 50,000 rounds period.
Figure 1: Average simple regret for each Best Arm Identification algorithm when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 2: Average simple regret for each Best Arm Identification algorithm when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
5.3.2 Ludii
The average regret of each best agent identification algorithm, for the Ludii game-agent dataset is shown in Figure 3. Similar to the GVGAI datasets, our proposed Optimistic-WS algorithm again provides a significant performance increase, achieving a simple regret that is 35.1% smaller than UCB-E and 55.8% smaller than uniform sampling when averaged across all 50,000 rounds.
One immediate observation is that the reduction in average simple regret for the Ludii dataset is less rapid compared to the GVGAI dataset. This could be due to several factors, but is most likely caused by the fact that the Ludii domain simply has more games compared to GVGAI, and hence more bandits to split pulls between. As a result, we would require a much larger number of rounds before our regret would begin to plateau, although we can see this beginning to occur for Optimistic-WS towards the 50,000 round mark.
Figure 3: Average simple regret for each Best Arm Identification algorithm when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
5.4 Discussion
Looking first at the GVGAI dataset results, see Figures 1 and 2, we can identify some general trends for each algorithm. Interestingly, the performance comparisons between the GVGAI (all levels) and GVGAI (level 1) datasets are remarkably similar, and so our observations can be jointly applied to both.
Optimistic-WS clearly performs best, achieving a lower average simple regret than all other algorithms for any number of rounds. After this, the second best performing approach is most often UCB-E, with the default UCB1 exploration value of $a=2\cdot\log(n)$. However, after this the comparison is not as clear cut, with different algorithms performing better after varying round amounts. Gape-E starts off well early on but performs worse as the number of rounds increases, eventually only beating Random and Uniform after 50,000 rounds. GapE-V by comparison starts off worse, but overtakes GapE around the 15,000 round mark. Both Successive Rejects and Sequential Halving are designed for a fixed total number of arm pulls, and so initially perform very poorly as the are not intended for early stopping. However, they eventually match the performance of other top performing models, such as UCB-E, once we reach defined 50,000 round end point. The Anytime Sequential Halving algorithm demonstrates its improved “stop anytime” functionality over regular Sequential Halving, achieving a much lower average simple regret in the earlier rounds and a near equal performance after 50,000 rounds. Lastly, it is unsurprising that Random and Uniform performed the worst overall, both of which are fairly naive (although still commonly used) evaluation approaches.
Looking next at the Ludii dataset results, see Figure 3, we can see that Optimistic-WS once again significantly outperforms all other approaches. UCB-E started off well, but eventually worsened and ended up matching the performance of GapE after 50,000 rounds). GapE-V also appeared to perform worse than its GapE counterpart, as did Anytime Sequential Halving. This apparent improvement in GapE compared to the GVGAI results could be due to the fixed 50,000 round limit, and that with additional rounds the relative performance of GapE may worsen (although this is unconfirmed). Lastly, the results for Successive Rejects and Sequential Halving are particularly odd, flattening well before the 50,000 round limit. The reason for this is largely due to their requirement of a fixed number of arm pulls per bandit, and is discussed in greater detail below.
One significant drawback of the Sequential Halving and Successive Reject algorithms is that they often do not use all of the pulls available to them (Audibert et al., 2010; Karnin et al., 2013). This is because both algorithms distribute pulls across several arm selection rounds ahead of time, requiring that the number of pulls each round be divisible between the remaining arms, which often results in some leftover pulls. This problem is particularly prevalent in situations where the number of available pulls is relatively low compared to the number of bandit arms. Our Ludii dataset presents just such a scenario, with 1085 games and a budget of 50,000 trials meaning that each game is assigned only 46 trials. Given that each game can have up to 29 potential agents that need evaluating, this very small trials-to-agents ratio severely limits the effectiveness of these approaches. Sequential Halving is most impacted by this, with only 26,834 ($\sim$54%) of its allocated trials being used, resulting in a flat regret line after this point. Successive Rejects is also affected by this same issue, using just 39,102 ($\sim$78%) of its available trials, but the fact that it discards its arms one at a time instead of removing half appears to reduce this impact. This issue is also much less prevalent for the GVGAI dataset, likely due to the reduced number of games this dataset has, leading to an increased number of trials per game-agent pair.
Overall, the improvements in best agent identification accuracy and efficiency offered by our proposed Optimistic-WS algorithm are clear across both domains, being able to to consistently outperform all other techniques for any number of rounds tested. The benefits of our Optimistic-WS algorithm appear most prominent in the early-mid round period, where it quickly achieved a very low simple regret compared to other techniques. As more rounds were performed we see that this gap begins to shrink, although the overall percentage improvement remains relatively stable. Given an infinite number of rounds, all techniques would (in theory) be able to obtain a perfect average simple regret score of zero, although such an assumption is clearly not feasible in practice. Demonstrating our Optimistic-WS algorithm’s improved performance over a wide range of possible rounds emphasises its practical application to estimating agent performance in a realistic experimental setting.
6 Conclusion
In this paper we have presented our novel Optimistic-WS algorithm for best arm identification in multi-armed bandits. This approach uses the Wilson score interval to estimate the regret change potential of each arm, and has been designed specifically for the task of identifying the best agents in a general game playing domain. Experiments conducted on two of the most popular general game systems (GVGAI and Ludii) reveal that our proposed approach significantly outperforms all previous state-of-the-art algorithms. We hope that this approach will be applied to enhance future agent evaluation procedures across a wide range of general game systems.
6.1 Future Work
There were several aspects that we did not consider for our presented Optimistic-WS algorithm that could be explored further. In our experiments, all trials were assumed to take an equal length of time to complete. However, this is often not the case, with certain games likely taking much longer to complete than others. Skilled agents may also be able to win easy games quicker than less skilled agents, or conversely take longer to lose at harder games. While we did not explore these aspects within this paper, it should be possible to create a “cost aware” version of our algorithm that considers the expected computational expense of performing a trial for a specific game-agent pair (based on prior experience). We also did not consider any domain specific measures that could be used to identify performance correlations between games/agents. If we are able to estimate that two (or more) games/agents have similar profiles, then results obtained from one trial could potentially be applied proportionately to several other related games/agents. Domain agnostic measures of game/agent distance could also be determined based on the similarity of prior results, although these may not always prove reliable.
Another suitable avenue for future work would be to identify additional multi-task, multi-agent domains for carrying out further evaluations. Our approach is general enough to work on datasets for any such domains, providing the requirement on reward limits being bounded between 0 and 1 is satisfied. While we selected the Wilson score confidence interval specifically for the domain of general game playing, it is also possible that other confidence intervals would be better suited for alternative domains with different reward distributions. Further evaluations will therefore be needed to accurately estimate the wider applicability of our approach.
Lastly, one of the immediate applications of our proposed best-agent-identification process would be to improve the performance of portfolio agents that leverage a suite of other pre-existing agents to make decisions. In their simplest form, portfolio agents are classification models trained to predict the best performing agent for a specific game. These agents are trained on labelled instances of prior games where the best performing agent is already known. Using our proposed best-agent-identification approach, the best performing agent in any given game can be estimated far more efficiently. This can potentially lead to a significant boost in the training speed and performance of portfolio agents for general game playing, as well as other domains where ground truth labels are expensive to compute.
A GVGAI (all levels) additional hyperparameter results
In this appendix we provide additional results for other algorithm hyperparameter values not shown in Figure 1
Figure 4: Average simple regret for each GapE hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 5: Average simple regret for each GapE-V hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 6: Average simple regret for each UCB-E hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 7: Average simple regret for each Optimistic-WS hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
B GVGAI (level 1) additional hyperparameter results
In this appendix we provide additional results for other algorithm hyperparameter values not shown in Figure 2
Figure 8: Average simple regret for each GapE hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 9: Average simple regret for each GapE-V hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 10: Average simple regret for each UCB-E hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 11: Average simple regret for each Optimistic-WS hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
C Ludii additional hyperparameter results
In this appendix we provide additional results for other algorithm hyperparameter values not shown in Figure 3
Figure 12: Average simple regret for each GapE hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 13: Average simple regret for each GapE-V hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 14: Average simple regret for each UCB-E hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 15: Average simple regret for each Optimistic-WS hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.